# Meta-learning-based percussion transcription and $t\bar{a}la$ identification from low-resource audio

### v2

@@ IV-A2 Step 2 - A): Model Agnostic Meta-learning (MAML)
Let, a task $T$ is sampled from a distribution $p(T)$ within the meta-training dataset $D_{train}$. Each task $T_{i}\sim p(T)$ is divided into a support set $T_{i}^{s}$ consisting of $s$ samples and a query set $T_{i}^{q}$ containing $q$ samples. The pre-trained base model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$, with parameters $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ explained in Section IV-A1, serves as the base-learner model. During meta-learning, we freeze the parameters of the convolutional layers and update the parameters of the recurrent and classifier layers. Thus, $\theta_{1}$ remains the same as in the pre-trained model while $\theta_{2}$ and $\theta_{3}$ become trainable parameters. We denote the combined parameters $\theta_{2}$ and $\theta_{3}$ as $\phi$, making $\phi$ the meta parameters of the base-learner model. The updated parameters of the base learner after $N$ steps for task $T_{i}$ are given by,
@@
where, $\alpha$ represents the base model’s learning rate, and $L_{T_{i}^{s}}(f_{[\theta_{1},\phi_{N-1}^{i}]})$ denotes the loss computed on the support set of task $T_{i}$ after $(N-1)$ update steps, as given in Equation (2). This process, known as inner loop optimization (ILO), involves updating the model $f_{[\theta_{1},\phi]}$ based on the support set. Once base learning is complete, the model $f_{[\theta_{1},\phi]}$ becomes $f_{[\theta_{1},\phi_{N}^{i}]}$. Subsequently, using these updated parameters $f_{[\theta_{1},\phi_{N}^{i}]}$, the query loss $L_{T_{i}^{q}}({\phi}_{N}^{i},T_{i}^{q})$ is computed on the query set $T_{i}^{q}$. The meta-parameters $\phi$ are then updated using the query loss. The process of updating meta-parameters over the batch of tasks is called outer loop optimization (OLO) and is given by,
@@
IV-A3 Step 2 - B): Transfer learning
We use pre-trained models from the mridangam dataset and fine-tune them with a small tabla stroke dataset to leverage their acoustic similarities. Transfer learning is applied to a pre-trained model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ with parameters $\theta_{1}$, $\theta_{2}$ and $\theta_{3}$, as described in Section IV-A1. During fine-tuning, $\theta_{1}$ remains fixed, while $\theta_{2}$ and $\theta_{3}$ are combined and updated as $\phi$. The parameters $\phi$ are optimized using the stochastic gradient descent algorithm Equation (1) and the weighted cross-entropy loss Equation (2).
IV-B $T\bar{a}la$ identification
$T\bar{a}las$ are primarily identified by their $th\bar{e}k\bar{a}$, which maintains a fixed stroke count as discussed in Section III. We utilize these sequence and ratio properties for $t\bar{a}la$ identification, defining two $t\bar{a}la$ identification methods (scores): ‘NW Matching Score’ and ‘Stroke Ratio Score,’ based on the Needleman-Wunsch algorithm [45] and Cosine Similarity, respectively.
@@ IV-B1 NW Matching Score
IV-B2 Stroke Ratio Score
@@
IV-C Automatic Drum Transcription (ADT)
This section presents our solution for ADT with minimal labelled data. We employ transfer learning and model-agnostic meta-learning (MAML) for ADT.
IV-C1 MAML for ADT
A similar MAML process, as explained in Section IV-A2, is adopted for ADT. This approach employs a CRNN model as the base learner for multi-label classification. The input feature matrix is a standardised log Mel-spectrogram, similar to those used in PTM1 and PTM2, with the target vector methodology following the approach described in [12, 13]. Unlike tabla stroke transcription, where the convolutional layer parameters are frozen during meta-training and meta-testing, all parameters $\theta_{1},\theta_{2},\theta_{3}$ are trainable in ADT. These parameters are optimised using the weighted binary cross-entropy loss and updated via the stochastic gradient descent algorithm.
IV-C2 Transfer Learning for ADT
A similar transfer learning approach, as detailed in Section IV-A3, is utilized for ADT. The model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ is initially trained on a source domain dataset and subsequently fine-tuned on a target domain dataset. Model parameters are optimized using the stochastic gradient descent algorithm with the weighted binary cross-entropy loss.
@@ V Experimental Setup
This study employs simple deep CNN and CRNN models for mridangam and tabla stroke transcription and the ADT task across all experiments. The architectures of the CNN and CRNN models are depicted in Fig. 3. Comprehensive details regarding the datasets utilized and the experimental settings for each method are outlined in the following sections.
@@
V-A Datasets
For ADT, we use the $D_{D1}$ : IDMT-SMT-Drums [46] (DTD type), $D_{D2}$: ENST-Drums minus-one [47] (DTP type), $D_{D3}$: MDB-Drums [48] (DTD and DTM type), and $D_{D4}$ : ADTOF [49] (DTM type).
For TST, we use the $D_{T1}$ : Tabla Solo Dataset [9], $D_{T2}$ : the 4way-tabla-ismir21-Dataset [5], and $D_{T3}$: the Hindustani Music Rhythm Dataset [43]. $D_{T1}$ and $D_{T2}$ are tabla-only datasets, while $D_{T3}$ includes tabla along with melodic instruments and vocals, forming a complete concert dataset. Additionally, we curated a Synthetic Mridangam Stroke Dataset $(D_{M})$, as explained below.
Synthetic Mridangam Stroke Dataset $(D_{M})$
We curated a dataset of 720 audio files, each with a duration of 10 seconds, by concatenating randomly selected mridangam strokes from the existing Mridangam Stroke Dataset [50]. This original dataset includes audio samples for ten distinct stroke labels across six different tonic pitches, recorded at a sampling rate of 44.1 kHz. The resulting dataset consists of 120 audio files for each tonic, amounting to a total audio duration of 120 minutes.
TABLE I: TST : Meta-train-test dataset
Train/Test Datasets Stroke Classes
Meta-train
$D_{T1}$
Da, Ki, Ge, Kda, Tit,
Dha, Dhe, Dhi, Re, Tin
Meta-test-1
$D_{T1}$ Ta, Na, Din, Dhet, Dhin
Meta-test-2
$D_{T2}$
B, D, RB, RT
Meta-test-3
$D_{T3}$
Dha, Dhin, Tin, Na, Tun, Kat,
Ta, Dhage, Tirkita, Dhi, Ti
TABLE II: ADT : Meta-train-test dataset
Train/Test Dataset Stroke classes
Meta-train $D_{D4}$ BD, SD, TT, HH, CY+RD
Meta-test-1 $D_{D1}$ KD, SD, HH
Meta-test-2 $D_{D2}$
BD, SD, CL, HH, BE,
TT, RD, CY
Meta-test-3 $D_{D3}$
KD, SD, CHH, HH, TT,
RD, CY, OT, TMB
V-B Tabla Stroke Transcription (TST)
V-B1 Step - 1: Pre-training on $D_{M}$
One-way Transcription Method
We employ 10 CNNs for 10 classes as discussed in V-A. Akshay et al. [51] categorized these classes into damped (D), resonant-treble (RT), resonant-bass (RB), and resonant-both (B). We use the same model architectures as Rohit et al. [5] for these ten classes as follows: for $Thi$, $Ta$, $Num$, $Tha-$D; for $Dhin$, $Cha$, $Bheem-$ RT; for $Thom-$RB; and for $Tham$, $Dheem-$B. To achieve tonic-independent inference, we applied 6-fold cross-validation, splitting the dataset into six folds by tonic. We trained models on five folds and tested on one, with each fold’s training data split 90:10 for training and validation. This process covered all six tonics and we averaged the evaluation scores.
The CNN processes log-Mel spectrograms from three channels with window durations of 23 ms, 46 ms, and 92 ms, all using a fixed hop size of 10 ms. Each channel’s input has dimensions of 128 Mel-frequency bands by 15 time frames. Training employs the Adam optimizer with a batch size 256 for up to 150 epochs. Early stopping with a patience of 10 epochs is used, saving the model with the lowest validation loss. During inference, a fixed-threshold peak picking algorithm [52] binarizes the output, identifying significant peaks to convert continuous values into binary decisions.
@@ V-B1 Step - 1: Pre-training on $D_{M}$
The datasets used in this work vary in audio length, so we segment them into nonoverlapping 5-second chunks for feature extraction, each labeled accordingly. The input to the CRNN is a standardized log Mel spectrogram, computed with a 46.4 ms window size and a 10 ms hop size, utilizing 128 Mel filter banks. Each frequency band of the Mel spectrogram is normalized to have zero mean and unit variance. This method assigns equal importance to the duration of the entire stroke rather than focusing solely on the onset frame.
Mridangam and tabla strokes are primarily characterized by the attack, decay, and sustain phases of the ADSR model, with most strokes having a fast-decaying exponential envelope [21]. Therefore, stroke onset is predominantly defined by the energetic attack-decay phase, while the sustain phase defines how the stroke ends. The last part of the release region contributes minimally to defining stroke onset and class. If an equal focus is given to the complete stroke duration during transcription, it can confuse the model since this phase is similar across all stroke classes. Hence, we transcribe this segment separately as ‘No-stroke’ using a threshold set at 3% of the stroke’s maximum amplitude. Frames from onset to threshold denote the stroke class, while frames from threshold to the subsequent stroke onset are labeled ‘No-stroke.’ We employ 6-fold cross-validation using six tonics. Training involves a batch size of 32 over 100 epochs, with a learning rate of 0.001. During inference, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class for evaluation.
@@
The CRNN model architecture and normalized log Mel spectrogram input to the CRNN are the same as explained in Section V-B1. The training targets include the ‘No-stroke’ class, as detailed in Section IV-A1. The model is trained with a batch size of 32 for 100 epochs, using a learning rate of 0.001. During inference, the model predicts stroke classes frame-by-frame for a given test log-Mel spectrogram. The onset of each stroke class corresponds to the frame in which the model predicts it. Then, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class.
V-B2 Step 2
Based on the pretraining results, one-way transcription performs poorly in complexity and time consumption, requiring a CNN for each stroke class. As a result, we will discontinue this approach and instead focus on the proposed methods. Specifically, we will employ the CRNN model for MAML-based domain adaptation and transfer learning-based domain adaptation for tabla strokes. Our methodology will be validated across three tabla datasets.
@@ V-B2 Step 2
We meta-train our models on ten stroke classes from $D_{T1}$ and conduct meta-testing in three scenarios: evaluating performance on five stroke classes from $D_{T1}$, testing with four stroke classes from $D_{T2}$, and assessing eleven stroke classes from $D_{T3}$. Additionally, we performed experiments using five, six, and eleven stroke classes from $D_{T3}$ to study the impact of class number in meta-testing. Detailed stroke class distributions for meta-training and meta-testing are in Table I. We randomly selected 12 minutes of audio from $D_{T1}$ for meta-training and used 5 minutes from $D_{T1}$ , 10 minutes from $D_{T2}$ , and 10 minutes from $D_{T3}$ for meta-testing. Furthermore, the model adapted during meta-testing on $D_{T2}$ was directly evaluated on a separate test partition from $D_{T2}$ provided by Rohit et al. [5], enabling comparison with the baseline approach.
@@
In the first scenario, we must partition the existing classes into two mutually exclusive sets. Due to the challenge of directly segmenting long audio based on stroke classes, we adopt an approach from Nolasco et al. [28] and Shi Bowen et al. [34]. Here, we preserve the original class label for desired strokes and classify non-target strokes as ’Others.’ As a result, there will be differences in the classes used between meta-training and meta-testing.
@@
Although some stroke types in $D_{T2}$ and $D_{T3}$ may appear similar, they are performed with different background instruments: $D_{T3}$ includes vocals and various instruments, while $D_{T2}$ is standalone. Therefore, we treat them as distinct stroke types. The ‘No-stroke’ class is also considered in the ‘PTM1’ and ‘PTM2’, discussed earlier in Sections IV-A1 and IV-A1, respectively. The MAML approach is applied to both methods with the same experimental settings, utilizing their respective input-output pairs. For task $T_{i}$, we randomly sample log Mel spectrograms with a batch size comprising 32 support samples and 8 query samples from the meta-training dataset. Models are trained with the Adam optimizer for 300 epochs, employing an inner loop optimization of $N=3$. The learning rates $\alpha$ and $\beta$ are fixed at 0.001.
Transfer Learning
We validate the transfer learning method using the same classes and samples from $D_{T1}$, $D_{T2}$, and $D_{T3}$ as in Section V-B2. Additionally, a fine-tuned model on $D_{T2}$ is tested on separate test data provided by Rohit et al. [5] in $D_{T2}$. Transfer learning is applied with the same experimental settings to both methods described in Sections IV-A1 and IV-A1, using their respective input-output pairs. The models are trained with a batch size of 32 for 100 epochs and a learning rate of 0.001.
V-C Automatic Drum Transcription (ADT)
V-C1 MAML for ADT
Meta-training is performed using five stroke classes from $D_{D4}$ while meta-testing is carried out in three distinct scenarios: (i) three stroke classes from $D_{D1}$, (ii) eight stroke classes from $D_{D2}$, and (iii) nine stroke classes from $D_{D3}$. The detailed distribution of stroke classes used for meta-training and meta-testing is provided in Table II. The complete $D_{D4}$ dataset is utilized for meta-training, and 5-minute excerpts from $D_{D1}$, $D_{D2}$, and $D_{D3}$ are used for meta-testing.

### v3

@@ III-A2 Step 2: Model Agnostic Meta-learning (MAML)
Let, a task $T$ is sampled from a distribution $p(T)$ within the meta-training dataset $D_{train}$. Each task $T_{i}\sim p(T)$ is divided into a support set $T_{i}^{s}$ consisting of $s$ samples and a query set $T_{i}^{q}$ containing $q$ samples. The pre-trained base model $f_{[\theta_{1},\theta_{2},\theta_{3}]}$, with parameters $f_{[\theta_{1},\theta_{2},\theta_{3}]}$ explained in Section III-A1, serves as the base-learner model. During meta-learning, we freeze the convolutional layers (pre-trained on mridangam) and update the recurrent and classifier layers. The convolutional layers capture low-level features shared between mridangam and tabla, preserving transferable features and reducing overfitting on low-resource tabla data. Thus, $\theta_{1}$ remains the same as in the pre-trained model while $\theta_{2}$ and $\theta_{3}$ become trainable parameters. We denote the combined parameters $\theta_{2}$ and $\theta_{3}$ as $\phi$, making $\phi$ the meta parameters of the base-learner model. The updated parameters of the base learner after $N$ steps for task $T_{i}$ are given by,
@@
where, $\alpha$ represents the base model’s learning rate, and $L_{T_{i}^{s}}(f_{[\theta_{1},\phi_{N-1}^{i}]})$ denotes the loss computed on the support set of task $T_{i}$ after $(N-1)$ update steps, as given in Equation (2). This process, known as inner loop optimization (ILO), involves updating the model $f_{[\theta_{1},\phi]}$ based on the support set. Once base learning is complete, the model $f_{[\theta_{1},\phi]}$ becomes $f_{[\theta_{1},\phi_{N}^{i}]}$. Subsequently, using these updated parameters $f_{[\theta_{1},\phi_{N}^{i}]}$, the loss $L_{T_{i}^{q}}(\phi_{N}^{i},T_{i}^{q})$ is computed on the query set $T_{i}^{q}$. The meta-parameters $\phi$ are then updated using this loss. The process of updating meta-parameters over the batch of tasks is called outer loop optimization (OLO) and is given by,
@@
III-B $T\bar{a}la$ identification
$T\bar{a}las$ are primarily identified by their $th\bar{e}k\bar{a}$, which maintains a fixed stroke count as discussed in Section II-C. We utilize these sequence and ratio properties for $t\bar{a}la$ identification, defining two $t\bar{a}la$ identification methods (scores): ‘NW Matching Score’ and ‘Stroke Ratio Score,’ based on the Needleman-Wunsch algorithm [39] and Cosine Similarity, respectively.
@@ III-B1 NW Matching Score
III-B2 Stroke Ratio Score
@@
III-C Automatic Drum Transcription (ADT)
This section presents our ADT solution under limited labelled data conditions. We adopt MAML for this task, following a process similar to that described in Section III-A2. A CRNN model serves as the base learner for multi-label classification. The input feature matrix is a standardized log Mel-spectrogram, consistent with those used in PTM1 and PTM2, and the target vector is constructed following the approach in [15, 16]. Unlike in TST, where pre-trained weights are used, no relevant pre-trained model exists for Western drum kits in ADT. Therefore, the entire model—including convolutional, recurrent, and linear layers—is trained from scratch, with all parameters $\theta_{1},\theta_{2},\theta_{3}$ kept trainable. where convolutional layer parameters are frozen during meta-training and meta-testing, all parameters $\theta_{1},\theta_{2},\theta_{3}$ are trainable in ADT. These parameters are optimized using a weighted binary cross-entropy loss and updated via stochastic gradient descent.
IV Datasets
We use diverse datasets for TST and ADT, including Only Drum/Tabla (DTD), Drums with Percussion (DTP), and Complete Concert Mixes (DTM), comprising drums/tabla alongside other percussion, melodic instruments, and vocals. Tables I and II summarize the datasets, detailing notations, durations, number of tracks/compositions, stroke classes, and data usage for meta-training/testing.
While some stroke or instrument classes (e.g., $Dha$, $Dhi$, snare, hi-hat) appear across datasets, they are treated as distinct during meta-learning due to differences in recording conditions, background instrumentation, and data distribution. Even with identical labels, contextual variations (e.g., presence of accompanying instruments or vocals) require treating them separately to ensure robust domain adaptation. Audio samples for meta-training and testing were randomly selected as per durations listed in Tables I and II.
TABLE I: Meta-Learning Dataset Configuration for Tabla Stroke Transcription (TST)
Dataset Notation Total Dur. (hr) Stroke Classes Used Meta Train-Test Data Usage & Evaluation
Synthetic Mridangam
Stroke Dataset
$D_{M}$
      2.0
(720 excerpts)
$Thi,Ta,Num,Tha,Dhin,Cha,$
$Bheem,Thom,Tham,Dheem$
Used in Pre-training stage
Tabla Solo
Dataset [25]
$D_{T1}$
      0.29
(38 composi-
      -tions)
$Da,Ki,Ge,Kda,Tit,Dha,Dhe,$
$Dhi,Re,Tin,Ta,Na,Din,Dhet,$
$Dhin$
Meta-Train: 12 min (0.2 hr) and 10 classes
($Da,Ki,Ge,Kda,Tit,Dha,Dhe,Dhi,Re,Tin$)
Meta-Test-1: 5 min (0.08 hr) adaptation on 5 new classes
($Ta,Na,Din,Dhet,Dhin$); Evaluation on 0.21 hr
4way-tabla-Dataset
ismir21 [6]
$D_{T2}$
      1.6
(38 tracks)
B, D, RB, RT
Meta-Test-2: 10 min (0.17 hr) adaptation on new classes;
Evaluation on a 20 min (0.33 hr) separate test set by [6]
Hindustani Music
Rhythm Dataset [34]
$D_{T3}$
        5.03
(151 excerpts)
$Dha,Dhim,Tin,Na,Tun,Kat,$
$Ta,Dhage,Tirikita,Dhi,Ti$
Meta-Test-3: 10 min (0.17 hr) adaptation on new classes;
Evaluation on 4.86 hr
TABLE II: Meta-Learning Dataset Configuration for Automatic Drum Transcription (ADT)
Dataset Notation ADT Task Type Total Dur. (hr) Stroke Classes Used Meta Train-Test Data Usage & Evaluation
ADTOF [40] $D_{D4}$ DTM
      114
(1739 tracks)
BD, SD, TT, HH, CY+RD Meta-Train: on full dataset (all classes seen)
IDMT-SMT-
Drums [41]
$D_{D1}$ DTD
     2.04
(104 tracks)
KD, SD, HH
Meta-Test-1: 5-min (0.083hr) adaptation
on new classes; Evaluation on 1.96 hr
ENST-Drums
minus-one [42]
$D_{D2}$ DTP
     1.01
(64 tracks)
BD, SD, CL, HH,
BE, TT, RD, CY
Meta-Test-2: 5-min (0.083 hr) adaptation
on new classes; Evaluation on 0.927 hr
MDB-Drums [43] $D_{D3}$ DTD, DTM
0.35
(23 tracks)
KD, SD, HH, HH-O,
TT, CY, RD, OT
Meta-Test-3: 5-min (0.083 hr) adaptation
on new classes; Evaluation on 0.267 hr
IV-A Synthetic Mridangam Stroke Dataset $(D_{M})$
The mridangam is a South Indian barrel-shaped percussion instrument with treble and bass heads, offering a wide tonal range. It is traditionally carved from jackfruit wood and fitted with animal skin membranes. We curated a dataset of 720 ten-second audio files by randomly concatenating strokes from the Mridangam Stroke Dataset [44], which contains 10 stroke classes across 6 tonic pitches, recorded at 44.1 kHz. The dataset includes 120 files per tonic, totalling 120 minutes of audio.
@@ V Experimental Setup
The experimental setup follows the same flow as shown in Fig.1 and consists of two major components: TST and ADT. Across all experiments, we employ simple deep CNN and CRNN models for mridangam and tabla stroke transcription, as well as for the ADT task. The architectures of the CNN and CRNN models are illustrated in Fig.3.
@@
V-A Tabla Stroke Transcription (TST)
V-A1 Step - 1: Pre-training on $D_{M}$
One-way Transcription Method (OTM)
We employ 10 CNNs for 10 classes as discussed in IV-A. Akshay et al. [45] categorized these classes into damped (D), resonant-treble (RT), resonant-bass (RB), and resonant-both (B). We use the same model architectures as Rohit et al. [6] for these ten classes as follows: for $Thi$, $Ta$, $Num$, $Tha-$D; for $Dhin$, $Cha$, $Bheem-$ RT; for $Thom-$RB; and for $Tham$, $Dheem-$B. To achieve tonic-independent inference, we applied 6-fold cross-validation, splitting the dataset into six folds by tonic. We trained models on five folds and tested on one, with each fold’s training data split 90:10 for training and validation. This process covered all six tonics, and we averaged the evaluation scores.
The CNN, trained on 3-channel log-Mel spectrogram inputs, uses the Adam optimizer with a batch size of 256 for up to 150 epochs. Early stopping with a patience of 10 epochs is applied, and the model with the lowest validation loss is retained. During inference, a fixed-threshold peak picking algorithm [46] is used to binarize the network output, identifying significant peaks and converting the continuous activations into binary decisions.
@@ V-A1 Step - 1: Pre-training on $D_{M}$
The datasets used in this work vary in audio length, so we segment them into non-overlapping 5-second chunks for feature extraction, each labelled accordingly. The input to the CRNN is a standardized log Mel spectrogram, computed with a 46.4 ms window size and a 10 ms hop size, utilizing 128 Mel filter banks. Each frequency band of the Mel spectrogram is normalized to have zero mean and unit variance. This method assigns equal importance to the duration of the entire stroke rather than focusing solely on the onset frame.
Mridangam and tabla strokes are primarily characterized by the attack, decay, and sustain phases of the ADSR model, with most strokes having a fast-decaying exponential envelope [21]. Therefore, stroke onset is predominantly defined by the energetic attack-decay phase, while the sustain phase defines how the stroke ends. The last part of the release region contributes limitedly to defining stroke onset and class. If an equal focus is given to the complete stroke duration during transcription, it can confuse the model since this phase is similar across all stroke classes. Hence, we transcribe this segment separately as ‘No-stroke’ using a threshold set at 3% of the stroke’s maximum amplitude. Frames from onset to threshold denote the stroke class, while frames from threshold to the subsequent stroke onset are labelled ‘No-stroke.’ We employ 6-fold cross-validation using six tonics. Training involves a batch size of 32 over 100 epochs, with a learning rate of 0.001. During inference, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class for evaluation.
@@
The CRNN model architecture and standardized log Mel spectrogram input to the CRNN are the same as explained in Section V-A1. The training targets include the ‘No-stroke’ class, as detailed in Section III-A1. The model is trained with a batch size of 32 for 100 epochs, using a learning rate of 0.001. During inference, the model predicts stroke classes frame-by-frame for a given test log-Mel spectrogram. The onset of each stroke class corresponds to the frame in which the model predicts it. Then, class-wise onsets are separated, and only ten stroke classes are considered by eliminating the ‘No-stroke’ class.
V-A2 Step 2
Based on the pretraining results, OTM performs poorly in complexity and time consumption, requiring a CNN for each stroke class. As a result, we will discontinue this approach and instead focus on the proposed methods. Specifically, we will employ the CRNN model for MAML-based domain adaptation for tabla strokes. Our methodology will be validated across three tabla datasets.
@@ V-A2 Step 2
We meta-train our models on ten stroke classes from $D_{T1}$ and evaluate in three meta-testing scenarios involving $D_{T1}$, $D_{T2}$, and $D_{T3}$. The corresponding stroke class distributions and audio durations for both meta-training and testing are detailed in Table I. All datasets come from different distributions and recording backgrounds. Additionally, the model adapted on $D_{T2}$ during meta-testing is directly evaluated on a separate test partition from $D_{T2}$ provided by Rohit et al. [6], allowing comparison with the baseline.
@@
In the first scenario, we must partition the existing classes into two mutually exclusive sets. Due to the challenge of directly segmenting long audio based on stroke classes, we adopt an approach from Nolasco et al. [26] and Shi Bowen et al. [31]. Here, we preserve the original class label for desired strokes and classify non-target strokes as ‘Others’. As a result, there will be differences in the classes used between meta-training and meta-testing. The ‘No-stroke’ class is also considered in the ‘PTM1’ and ‘PTM2’, discussed earlier in Sections III-A1,V-A1, and III-A1.
@@
The MAML approach is applied to both methods with the same experimental settings, utilizing their respective input-output pairs. For task $T_{i}$, we randomly sample log Mel spectrograms with a batch size comprising 32 support samples and 8 query samples from the meta-training dataset. Models are trained with the Adam optimizer for 300 epochs, employing an inner loop optimization of $N=3$. The learning rates $\alpha$ and $\beta$ are fixed at 0.001.
To evaluate the effectiveness of our proposed method against existing approaches, we conducted comparative experiments using a standard supervised training setup. Specifically, we trained and tested the same CRNN model using the PTM1 and PTM2 paradigms, with limited training data matching the meta-adaptation setup. We also evaluated a transfer learning baseline, where the model was fine-tuned and tested using the same adaptation and evaluation data. All experiments—supervised, transfer learning, and meta-learning—were performed on identical train-test splits to ensure fair comparison. The models were trained with a batch size of 32 for 100 epochs using a learning rate of 0.001.
V-A3 Baseline Models for TST Comparison
To evaluate the effectiveness of our proposed method, we conducted comparative experiments using a standard supervised setup. We trained and tested the same CRNN model with PTM1 and PTM2, using limited data matching the meta-adaptation setup. We also evaluated transfer learning, where the model was fine-tuned and tested with the same data. All experiments (supervised, transfer learning, and meta-learning) used identical train-test splits for a fair comparison. The models were trained with a batch size of 32 for 100 epochs at a learning rate of 0.001.
V-B Automatic Drum Transcription (ADT)
V-B1 MAML for ADT